The table shows the number of tulip bulbs Chloe and Grady planted. If they each planted the same number of bulbs, how many did each plant? Chloe= 3(t+1) Grady= 3(2t-3)
Answer by checkley79(3341)

3(t+1)=3(2t-3)
3t+3=6t-9
3t-6t=-9-3
-3t=-12
t=-12/-3
t=4 bulbs ans.
Proof:
3(4+1)=3(2*4-3)
3*5=3(8-3)
15=3*5
15=15